LUCIFER Com 3, Night 8
UT 2009 January 17

Observers: N.Ageorges, W.Seifert, M.Juette, V.Knierim, J.Hill (@LBT)
Telescope Operator: D.Gonzalez-Huerta
SW Support: P.Grenz in Tucson until 1 AM
Telescope Support: J.Rousseau

Summary:

Sky was photometric in the early evening, but broken clouds moved in by midnight. Seeing improved from 0.7 arcsec early to 0.4 arcsec by 10 PM. Closed by clouds at 1:15 AM, a reopened in partly cloudy conditions at 4:30 AM.

Lost half an hour at the beginning because we couldn't find stars. Was OK after we reset encoders. Don't understand what was wrong.

Lost another half hour to various TCS hanging problems.

We refined the measurement of LEFTZEROPOINT from last week.

Tested some "click-to-offset" LUCIFER software for spectroscopic setup. Took about an hour of spectroscopic data.

Preparations:

Day crew repaired the broken oil level sensor in HBS. (Which is why the WinCC forces were off - they had the power to HBS off while working on it.) This value for the oil level should no longer need to be forced.

Tom S. updated the fiducial values for the elevation encoders to try to fix EL encoder 0. See his telescopework email for details.

JMH killed a stray LBTB firefox on lbtmu03 that was using 99% of the CPU.

~16:50 LUCIFER taking calibrations in the dark dome.

17:05 stopped and restarted PSFGUI (memory), GCSL (memory), IIF(not talking to LUCIFER) plus PSFL and PCS incidentally.

17:16 Tripped the MCR relay for telescope. Why? Lost 5 min.

17:20 copied oacontrol.conf.sfpagw9 into /etc/oacontrol/oacontrol.conf to use Jesper's newest rotational tweak of the transformations. Killed and restarted oacserver.

Details (times in MST below)

Open

17:22 Open dome facing East, T=2.4 degC, D=-23 degC, wind calm, photometric except the distant southwest. Still using only Left Shutter Door. Vent Doors open.

17:28 HBS shuts off because WInCC forces were off. Took 10 minutes to recover when it should have taken 5 because somebody removed the piece of paper from the computer room with the instructions.

Flats

17:40 Slewing to zenith for LUCIFER sky flats.

17:50 Reset the Yglobal offsets on both M1 and M2 to be +2.8 mm above nominal. Then send -5000 nm Z7 toi M1 to cancel the initial coma.
M1: Xglobal=0 Yglobal=+2.8 Zglobal=0 RX=0 RY=0
M2: Xglobal=-7.010 Yglobal=+1.662 Zglobal=0 RX=114.28 RY=181.80

17:55 Kill another large PSFGUI and LSSGUI. Now lbtmu01 is responsive again.

18:20 Testing some LUCIFER software mods.

Stars or not?

18:54 preset TRACK

19:02 Adjust M1 Zglobal to +0.2 mm.

19:05 Making some offset tests

19:10 preset TRACK PARALLACTIC - did not find the expected pointing star.

19:20 Adjust M1 Zglobal to +0.4 mm

19:25 preset TRACK PARALLACTIC to another star. Didn't find that one either, but we have background stars.

19:27 Stop to reset telescope encoders. And we talked to Tom to confirm that the fiducial changes he made this morning should not have changed pointing. He also confirmed that he reset the EL encoders himself at zenith this morning. Our first interesting result was the EL encoder 0 seemed to succeed in the reset process where it was predicted to fail. It was left armed at the end of the process. It took us three attempts to get the AZ encoders to reset. We spent half an hour doing the recovery procedure, but we have not learned what the actual problem was.

19:53 preset TRACK PARALLACTIC - succeeded and we see the star. Adjust IE=39.0 CA=8.6

19:58 Testing Click-to-offset feature, and the star moved to where Marcus clicked.

20:04 preset ACTIVE off-axis - WFS seems to be converging slowly (6 iterations?). Seeing is 0.7 arcsec FWHM. Note that once again we find ourselves near the -Y limit at -2.49. We seem to be losing 100 microns per day? Or is it a temperature effect (presently 2.9 degC)?

20:18 FCin=1 for Z11/Z22

Spectroscopy Observing

20:46 OSS died. Paul restarts it.

20:50 Guiding images stop appearing.
Jan 16 20:51:26 lbtmu107 LBT_left_GCS: AzCam: IST reported error while receiving image.
Jan 16 20:51:26 lbtmu107 LBT_left_GCS: error: GuidingThread: guide image didn't arrive within timeout of 4388 ms.

Kill and restart GCSL.

20:54 resend preset ACTIVE

21:05 Seeing has improved below 0.5 arcsec.

21:20 Seeing has improved below 0.45 arcsec (0.43x0.39), and the local seeing perturbations to the WFS have settled down.

21:33 Lost guide star to calibration unit.

21:40 Problems finding star again.

Stop and restart PCS (nothing was wrong with PCS), but it isn't making syslog messages (Michele must have polys in quiet mode).

21:48 preset TRACK, but still not stars. It turned out to be the calibration unit still in place - observer error.

21:57 made a nice pupil image of sky and swing arms with pupil viewer in

22:02 Adjust IE=33.0 CA=5.0

Checking TRACK mode offsets

22:04 preset ACTIVE

22:15 Seeing has improved below 0.4 arcsec. Polaris monitor reports 1.2 arcsec.

22:24 preset ACTIVE - but WFS is saturated

22:30 Changed Filter to blue with "setfilter -f 0". The corresponding GCS "setFilter 1" command failed. (This would not help the WFS exposure time, since the filters are only in the guide arm. - Jesper)

22:38 Changed minimum guide exposure to 200 ms and minimum WFS exposure to 5000 ms in AIP_L.cfg. (These should not be permanent pending more data analysis by Doug.)

22:40 preset ACTIVE - now running 6752 ms WFS exposures with 29.7 sec cycle times. The preset put the filter back to r'. This data is also quite close to what Doug wanted for his fast WFS experiment. In general things worked OK with these short exposures.

22:56 preset TRACK - checking LEFTZEROPOINT with TRACK mode offsets

23:10 Change LEFTZEROPOINT in PCSInstrument.conf back to the previous value of -187.433. It was last changed by Dave at 00:10 on Jan 10 to be -187.726.

23:11 Stop PCS and restart.

23:12 preset ACTIVE - failed no star

22:02 Adjust IE=33.0 CA=5.0 since stopping PCS cleared them out.

23:14 preset ACTIVE

23:19 Change LEFTZEROPOINT in PCSInstrument.conf to -187.528 . Stop PCS and restart. Adjust IE=33.0 CA=5.0

Jan 16 23:20:48 lbtmu107 LBT_left_GCS: WFS: error: triggerWFS: WFS image didn't arrive within timeout of 13086 ms.

23:23 Kill and restart GCSL

23:24 preset ACTIVE - still with the bright guide star

23:29 preset TRACK

Clouds have appeared for the first time in many days.

23:32 Change LEFTZEROPOINT in PCSInstrument.conf to -187.433 . Stop PCS and restart. Adjust IE=33.0 CA=5.0

23:33 preset TRACK

23:37 preset ACTIVE

23:51 Change LEFTZEROPOINT in PCSInstrument.conf to -187.608 . Stop PCS and restart. Adjust IE=33.0 CA=5.0

23:52 preset ACTIVE

00:02 All three of our GCSGUIs have frozen. After we restart them, they freeze again. Based on experience of a few days ago, this could be a precursor of GCS hanging.

00:03 preset ACTIVE

00:11 Change LEFTZEROPOINT in PCSInstrument.conf to -187.565 . Stop PCS and restart. Adjust IE=33.0 CA=5.0

00:12 preset ACTIVE

00:17 Change LEFTZEROPOINT in PCSInstrument.conf to -187.651 . Stop PCS and restart. Adjust IE=33.0 CA=5.0

00:17 preset ACTIVE

00:20:33 GCS freezes on TRACK mode preset

00:23 Paul tries to stop GCSL, and it stops without a kill.

00:25 Paul restarts GCSL

00:25 preset TRACK

-187.651 degrees is the adopted value of LEFTZEROPOINT.

This marks the end of the short exposure WFS images for now.

Checking Pupil Alignment

00:34 Offset with 0, 0, 180deg

Compare LUCI images 20090116 0134 and 0135. It appears that the pupil misalignment is about half from the chief ray of the telescope and half from internal LUCIFER. A small rotation of the M3 mirror seems to be needed.

00:46 preset ACTIVE - couldn't find star - clouds?.

01:00 Priority shifts from observing, to barbequeing steaks on the patio.

01:15 The broken clouds are getting lower and heavier, so we close the dome.

Closed dome tests

01:20 Closed dome tracking for testing LUCIFER internal flexure compensation.

John went to bed around 3AM - a classic strategy for getting the clouds to clear. (His cold really feels much better after 10 hours of sleep.) Thus, this log ends here.

Re-open

The clouds thinned out, so everybody else reopened and observed from 04:30 until dawn. The clouds were nearly all gone by morning twilight.

Doug moved the guider images from 090116, about 3-5am into /Repository/AGW_Data/090116g so they don't get erased on the next reboot of lbtmu107.

-- JohnHill - 17 Jan 2009
Topic revision: r5 - 18 Jan 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback